Suppose the market has 70% chance of being favourable and 30% chance of being unfoavrable. A favourable market will yield a profit of N$300,000, while an unfavourable market will yield a profit of N$20,000. What is the expected monetary value in this situation?


1
Expert's answer
2022-04-20T14:10:04-0400

Let X - the random variable of the profit. It may take on the values 300000 and 20000 with the probabilities 0.7 and 0.3 respectively. The expected value of X is

"\mu=x_1\cdot P(X=x_1)+x_2\cdot P(X=x_2)=\\
=300000\cdot0.7+20000\cdot0.3=216000."